OpenStudy (anonymous):

For the function f(x)=4/x+3 average rate of change of f from 1 to x. How to even start with this?

OpenStudy (aum):

Average rate of change of function from a to b is: { f(b) - f(a) } / (b - a) Average rate of change of function from 1 to x is: { f(x) - f(1) } / (x - 1) Find f(1) and substitute into the above formula and simplify.

OpenStudy (anonymous):

So f(1)= 4/ (1) +3 4/4=1?

OpenStudy (aum):

Yes. Also substitute f(x) = 4 / (x+3) in the formula and simplify.

OpenStudy (anonymous):

Example? 4/ (x+3) x+3?

OpenStudy (aum):

Average rate of change of function from 1 to x is: { f(x) - f(1) } / (x - 1) f(x) = 4 / (x+3) f(1) = 4 / (1+3) = 4/4 = 1 { f(x) - f(1) } / (x - 1) = { 4 / (x+3) - 1 } / (x - 1) = { 4 - x - 3 } / { (x+3)(x-1) } = (1-x) / { (x+3)(x-1) } = - (x-1) / { (x+3)(x-1) } = -1 / (x+3)

OpenStudy (anonymous):

Thanks!

OpenStudy (aum):

You are welcome.